Operations Report – May 4th

Crew 261 Operations Report 04-05-2023 Update #1
SOL: 4
Name of person filing report: Julien Villa-Massone
Non-nominal systems:
· Remote observatory
· Helmet and headset
Notes on non-nominal systems:
· Remote observatory: N/A
· Helmet and headset: See detail below
Spirit rover used: Yes
Hours: (before EVA) 220.3
Beginning charge: (Before EVA) 100
Ending charge: (On return from EVA, before recharging) Not available due to priority given to simulated medical emergency. During second EVA (that did not use rovers), SOC of 81 was noted, while charging.
Currently charging: Yes
Opportunity rover used: No
Hours: 123.0
Beginning charge: 100
Ending charge: N/A
Currently charging: Yes
Curiosity rover used: No
Hours: 232.9
Beginning charge: 100
Ending charge: N/A
Currently charging: Yes
Perseverance rover used: Yes
Hours: 263.1
Beginning charge: 100
Ending charge: Not available due to priority given to simulated medical emergency. During second EVA, SOC of 100 was noted, while charging
Currently charging: Yes
General notes on rovers: Nominal operation today
Summary of Hab operations: Nominal day
WATER USE: rate of 30 gallons / sol. Trend of 28 Gallons/sol since mission start. We are now on target to finish the mission 37 gallons short (or run out 1.2 days early). This means, if we continue with this trend, the tank will be at 118 gallons when the mission ends.
Water (static tank): 349 gallons remaining
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no

Summary of internet: Nominal
Summary of suits and radios: Suits and radios:
One helmet is currently unusable due to a few screws being detached, no longer fixing the transparent part of the helmet to the metallic collar. The screws cannot readily be attached again because the metallic collar is damaged where the screws were located, and repair will be required. Crew Engineer does not believe a quick fix is possible, if the goal is not to break again during an EVA. It has been separated from the rest and labeled.
The headset microphone for the CDR’s radio (#1) operated unreliably. There were occasions when the CDR would press to talk and we would only hear static but not his voice. The microphone mounted on the radio functioned properly, so it is a headset / connector issue. Crew Engineer to investigate further when time allows tomorrow.
Summary of GreenHab operations: Nominal. Doors closed early due to wind.
WATER USE: 6 gal at 06:30 + 5 gal at 18:45 + 3 gal for research = 14 gal total
Heater: Off
Supplemental light: Off
Harvest: None
Summary of ScienceDome operations: Nominal. Early growth biostimulation setup.
Dual split: Off
POWER report: Nominal operation, except for a breaker disconnected between a PV string and an MPPT. The same as 2 days ago. Mission Support was alerted and immediately handled this issue. The disconnection happened between 9:27 (time of a photo) and 15:28 (time of discovery).
Summary of RAM operations: Crew Robotics Engineer continued assembly and adjustments of the robot prior to and following a second EVA near the base for testing.
Summary of any observatory issues: Nothing to report
Summary of health and safety issues: Nothing to report
Questions, concerns and requests to Mission Support: The entire crew joins me in thanking Mission Support and the entire Mars Society organization for making such an amazing experience possible.

Copyright © The Mars Society. All rights reserved. | Main Site